OpenStudy (anonymous):

What to do to solve this problem? find the difference in quotient of F; [f(x+h)-f(x)]/h; h can't equal 0. f(x)=x^2-x+4

OpenStudy (chaise):

Find the derivative by first principle?

OpenStudy (anonymous):

all i know is that it equals 8x+4h+5 but i'm not sure how it's solved =[

OpenStudy (chaise):

That's the formula for derivatives. Do you want me to do this for you and we'll see what happens? Are you sure it's "As h approaches 0", or is it definantly "h can't equal 0"

OpenStudy (anonymous):

easy, so, F(x) = x^2 - x + 4 F(x+h) = (x+h)^2 -(x+h) +4 Now, the formula states, [f(x+h)-f(x)]/h so, =(((x+h)^2 -(x+h) +4) - (x^2 - x + 4)))/h Expand =((x^2+2xh + h^2 - x + h + 4) - (x^2 - x + 4))/h x^2 is gone x is gone 4 is gone This leave =(2xh + h^2 + h)/h factor out h =h(2x+h+1)/h h cancels out =2x+h+1 As h approaches 0 =2x+0+1 =2x+1

OpenStudy (anonymous):

it says h≠0. I'm just not sure what the equation give is for or how its used to solve the function. Any ideas?

OpenStudy (anonymous):

That's the derivative by first principle

OpenStudy (chaise):

Should be 2x-1 rather than 2x+1. Don't ask me where Spykez went wrong.

OpenStudy (anonymous):

i see where i went wrong, it's -x-h, not -x+h

OpenStudy (anonymous):

so what is h or where did it come from?

OpenStudy (anonymous):

the derivative is a formula for the slope of the line tangent to the curve. your curve is \[y=x^2-x+4\] suppose you let x = 2 you get y = 6 so the point (2,6) is on the curve. then if you let x = 3 you get 10 so the point (3,10) is on the curve and the slope between the two point is 4 in my example x = 2 and h = 1 because 2+1=3 so i computed for the slope \[\frac{f(2+1)-f(2)}{1}\] which looks just like \[\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

now if you want the slope of the line between (2,6) and (2.1,f(2.1)) you get it by computing \[\frac{f(2.1)-f(2)}{.1}\] in this case x = 2 and h = .1

OpenStudy (anonymous):

by the way people who answered the question were anticipating the derivative, but your question says nothing about letting h go to zero, so it is just pure algebra. your answer would be \[2x-1+h\]